OpenStudy (anonymous):

How much of the circle is shaded? Write your answer as a fraction in simplest form. (circle below)

OpenStudy (anonymous):



OpenStudy (texaschic101):

1/4 + 1/7 = 7/28 + 4/28 = 11/28 leaving 17/28 which is 61% 11/28 = 39% (not shaded) 17/28 = 61% (shaded) -------------------add 28/28 = 100%

OpenStudy (texaschic101):

you might want to get a second opinion on this

OpenStudy (anonymous):

1/4= 7/28 1/7= 4/28 7+4= 11 28-11=17 answer= 17/28

OpenStudy (texaschic101):

that looks easier...lol...however, it is the same answer....thats good

OpenStudy (anonymous):

haha yeup!